Consider the following data (in millions) from Trident Financial Inc., which has two main divisions, mortgage loans and consumer loans:

Mortgage Loans Consumer Loans
Average total assets $ 2,700 $ 17,250
Operating income $ 540 $ 2,070
Return on investment (ROI) 20.00 % 12.00 %

Required:

1. Based on ROI, which division is more successful?

  • Mortgage loan or Consumer loans

2. Trident uses residual income (RI) as a measure of the financial performance of its divisions. What is the RI for each division if the minimum desired rate of return is (a) 10%, (b) 15%, and (c) 20%? (Leave no cells blank. Enter "0" wherever required. Negative amounts should be indicated by a minus sign. Enter your answers in millions of dollars, rounded to nearest whole number.)
Answer #1

1. As ROI is higher,

Mortgage Loans division is more successful

2. Residual income = Net operating income - Required return

Mortgage loans Consumer loans
10% 540 - (2700*10%) = 270 2070 - (17250*10%) = 345
15% 540 - (2700*15%) = 135 2070 - (17250*15%) = 518
20% 540 - (2700*20%) = 0 2070 - (17250*20%) = -1380